### v2

# Beyond Linear Probes: Dynamic Safety Monitoring for Language Models

Abstract
Monitoring large language models’ (LLMs) activations is an effective way to detect harmful requests before they lead to unsafe outputs. However, traditional safety monitors often require the same amount of compute for every query. This creates a trade-off: expensive monitors waste resources on easy inputs, while cheap ones risk missing subtle cases. We argue that safety monitors should be flexible–costs should rise only when inputs are difficult to assess, or when more compute is available. To achieve this, we introduce Truncated Polynomial Classifiers (TPCs), a natural extension of linear probes for dynamic activation monitoring. Our key insight is that polynomials can be trained and evaluated progressively, term-by-term. At test-time, one can early-stop for lightweight monitoring, or use more terms for stronger guardrails when needed. TPCs provide two modes of use. First, as a safety dial: by evaluating more terms, developers and regulators can “buy” stronger guardrails from the same model. Second, as an adaptive cascade: clear cases exit early after low-order checks, and higher-order guardrails are evaluated only for ambiguous inputs, reducing overall monitoring costs. On two large-scale safety datasets (WildGuardMix and BeaverTails), for 4 models with up to 30B parameters, we show that TPCs compete with or outperform MLP-based probe baselines of the same size, all the while being more interpretable than their black-box counterparts. Our code is available at https://github.com/james-oldfield/tpc.
1 Introduction
Recent years have seen a marked improvement in the capabilities of large language models (LLMs). Specifically, the emerging paradigm of test-time compute has led to numerous breakthroughs in reasoning (Guo et al., 2025; Wei et al., 2022), mathematics (Kojima et al., 2022), and coding (Wang et al., 2024) tasks alike. The central idea is simple: rather than allocating extra resources to pre-training, compute is spent dynamically during inference instead–providing an additional axis along which to scale model capabilities. Beyond maximizing performance at all costs, a key strength of this modern approach lies in the flexibility it affords. Compute can be spent only when the problem demands it, or when budget permits.
However, despite widespread benefits to model capabilities, dynamic computation (Han et al., 2021) for AI safety remains nascent. This is particularly true in the domain of LLM safety monitors, trained to detect harmful requests (Han et al., 2024), or problematic model behavior (Goldowsky-Dill et al., 2025; MacDiarmid et al., 2024; Chaudhary & Barez, 2025). Popular monitoring techniques include LLM-as-judges of natural language on the one hand (Inan et al., 2023; Zeng et al., 2025), and cheap linear probes in activation-space on the other (Alain & Bengio, 2017). In both cases, we argue that current approaches are inflexible. Considering that most requests are benign, dedicated LLMs have an excessively large minimum cost as always-on monitors, while activation probes provide only the most basic, static guardrails. Whilst recent work proposes to chain these two existing approaches (McKenzie et al., 2025; Cunningham et al., 2025), requiring large external LLMs that need finetuning/prompting limits their flexibility. In contrast, a single activation monitor that scales dynamically offers two key benefits:
  1. 1.
    One model, multiple safety budgets: monitors that scale with compute offer a flexible way to navigate the cost-accuracy trade-off. A single model can be evaluated with varying amounts of compute to meet different safety requirements.
  2. 2.
    Not all queries require strong monitors: dynamic models can adapt their defense to each input–keeping monitoring costs low for easy cases, only evaluating stronger guardrails when difficulty necessitates.
Figure 1: Dynamic activation monitoring with truncated polynomial classifiers: ① We train an order-NN polynomial in the LLMs’ activations $\bm{z}\in\mathbb{R}^{D}$ as a binary classifier of harmful prompts. ② At test-time, any number of $n\leq N$ terms can be evaluated to fit a variety of compute budgets; higher-order terms providing stronger guardrails only when necessary.
In this paper, we propose truncated polynomial classifiers (TPCs) to achieve these two properties–refining linear probes’ decision boundary by modelling rich higher-order interactions. Specifically, we show how a degree-$N$ polynomial can be trained and evaluated progressively, yielding $N$ nested submodels. Once trained, a single TPC provides dynamic defense across a range of compute budgets, through truncated evaluation. Evaluating higher-order terms provides stronger guardrails when needed, naturally generalizing the familiar linear probe (as illustrated in Fig. 1).
The prevailing ‘linear representation hypothesis’ holds that many high-level concepts are represented as one-dimensional subspaces in activation-space (Park et al., 2023). However, there is increasing evidence that not all features exhibit such simple linear structure (Engels et al., 2025; Smith, 2024). To build robust, general-purpose monitors, more powerful alternatives to linear probes are therefore needed. However, unlike existing non-linear models (e.g., MLPs), TPCs’ form remains intrinsically interpretable (Dubey et al., 2022) at moderate degrees. Higher-order polynomials model multiplicative interactions between LLM neurons (Jayakumar et al., 2020), explicitly modeling how they jointly contribute to the safety classification. As a result, TPCs give built-in classifier attribution to specific combinations of neurons (Pearce et al., 2025), providing transparency into the classifiers’ decisions in addition to strong monitoring performance.
Exhaustive experiments across $4$ LLMs (with up to 3030B parameters), multiple layers, and $2$ large-scale safety datasets show TPCs compete with or outperform MLP-based probes when parameter-matched, all the while offering built-in feature attribution. On certain LLMs, we find TPCs evaluated at a fixed-order bring up to $10\%$ improvement in accuracy over linear probes (for classifying particular categories of harmful prompts), and up to $6\%$ over MLP baselines. Furthermore, we show cascaded TPC evaluation yields performance on par with the full polynomial–yet requiring only slightly more net parameters than the linear probe. Our contributions are summarized below:
  • We propose truncated polynomial classifiers and a progressive training scheme to scale LLM safety monitoring with inference-time compute–extending the familiar linear probe with rich non-linear interactions.
  • We demonstrate two complementary evaluation modes of TPCs; user-driven evaluation to meet safety budgets and token-driven compute, conditional on input ambiguity.
  • Across $16$ layers in $4$ LLMs, we show that TPCs compete with or outperform (parameter-matched) MLP baselines monitoring for harmful requests on $2$ large-scale datasets–all the while offering built-in feature attribution.
2 Related work
External LLMs as monitors
Safety training of LLMs is a standard technique for preventing models from responding to problematic requests, either via post-training (Ouyang et al., 2022; Haider et al., 2024; Yuan et al., 2025; Bai et al., 2022) or during pre-training itself (O’Brien et al., 2025; Chen et al., ). Unfortunately, many models still remain vulnerable to attacks and jailbreaks (Hughes et al., 2024; Anil et al., 2024), underscoring the importance of post-training safety guardrails. One common strategy to achieve this is to use standalone LLMs trained as safety classifiers (Weng et al., ; Inan et al., 2023; Han et al., 2024; Zeng et al., 2025), leveraging LLMs’ ability to generalize to identifying novel categories of harmful inputs/responses (Inan et al., 2023). However, whilst LLMs-as-monitors are powerful, they bring significant computational cost on top of every request (Li et al., 2025), which can be prohibitively expensive for always-on monitoring.
Feature probing
One compelling alternative to using external LLMs to monitor natural language requests/responses is classifiers on LLMs’ internal activations–motivated by the idea that high-level concepts are often encoded in the intermediate representations (Park et al., 2023; Mikolov et al., 2013). In particular, Alain & Bengio (2017) proposes the use of simple ‘linear probes’ to assess the linear separability of features within a deep feature space. Further studies explore more complex model forms (White et al., 2021), with Pimentel et al. (2020a) suggesting probe accuracy should be considered as a function of complexity. Moreover, further work explores the extent to which the accuracy of linear probes provides evidence of target concepts being well-represented in the embeddings (Hewitt & Liang, 2019; Saphra & Lopez, 2019; Pimentel et al., 2020b). Relevant to this paper, White et al. (2021) proposes the use of a polynomial kernel as a non-linear probe. Through the use of the kernel trick, however, there is no explicit computation of terms of increasing degree that facilitate the progressive evaluation proposed in this work. Simple linear probes provide a powerful way to monitor for a range of concepts related to the safety of LLMs–such as catching sleeper agents (MacDiarmid et al., 2024), monitoring for factual awareness (Tamoyan et al., 2025), or truthfulness (Burns et al., 2023). Moving beyond probes on the activations directly, recent work (Bricken et al., 2024) explores probing Sparse Autoencoder features (Huben et al., 2024), and/or activations from prompting instructions to improve classification (Tillman & Mossing, 2025).
Cascades & ensembles
Combining or learning multiple submodels is a powerful way to improve upon single models. Multiple classifiers used in a cascade (or networks with early exits) bring robustness and/or speed in computer vision (Viola & Jones, 2004; Bourdev & Brandt, 2005; Romdhani et al., 2001), machine learning (Grubb & Bagnell, 2012; Xu et al., 2012), and deep neural networks (Teerapittayanon et al., 2016; Raposo et al., 2024; Yue et al., 2024) alike. Recent work similarly explores the combination of multiple models for LLM monitoring (McKenzie et al., 2025; Hua et al., 2025; Cunningham et al., 2025). Concretely, McKenzie et al. (2025); Cunningham et al. (2025) both show large computational savings using activation probes as a first line of two-stage defense–routing inputs to external LLM-as-monitors when uncertain. Whilst well-positioned to benefit from future LLM advances, both McKenzie et al. (2025); Cunningham et al. (2025) require additional LLM fine-tuning or prompting, and calls to extra LLMs during inference time. Instead, TPCs learn dynamic $N$-layer defense from neuron interactions, directly in the original LLMs’ activation space–offering built-in neuron attribution. We view these methods as complementary; in principle, a cascade of depth $N+1$ could combine TPCs with an LLM-as-monitor final layer for additional defense.
Polynomial neural networks
There has been a surge of interest in learning higher-order polynomials due to their attractive theoretical properties (Stone, 1948), finding application in generative (Chrysos et al., 2020; 2022) and discriminative models alike (Gupta et al., 2024; Babiloni et al., 2023; Chrysos et al., 2023). Through modeling higher-order interactions (Jayakumar et al., 2020; Novikov et al., 2016), recent work has advocated for variants of polynomials as inherently interpretable architectures (Pearce et al., 2025; Dubey et al., 2022). Our paper builds off this literature, proposing truncated evaluation as a mechanism for turning polynomials into dynamic models.
3 Methodology
We now introduce the truncated polynomial safety classifier. We first recall the preliminaries in Section 3.1. We then describe in Section 3.2 how polynomials extend probes for dynamic evaluation–detailing the proposed progressive training in Section 3.2.1 and cascading defense in Section 3.2.2.
3.1 Preliminaries
Notation
We denote matrices (vectors) using uppercase (lowercase) bold letters, e.g., $\bm{X}$ ($\bm{x}$), scalars in lowercase, e.g., $x$, and higher-order tensors in calligraphic letters, e.g., $\bm{\mathcal{{X}}}$. An element of an $N^{\text{th}}$-order tensor $\bm{\mathcal{{X}}}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{N}}$ is indexed by $N$ indices, written as $\bm{\mathcal{{X}}}(i_{1},i_{2},\ldots,i_{N})\triangleq x_{i_{1}i_{2}\ldots i_{N}}\in\mathbb{R}$. We use square brackets to group weights related to the $k$-th order term in a polynomial, e.g., $w^{[k]}$. Finally, for multiple summations sharing the same upper-bound, we use the shorthand $\sum_{d_{1},d_{2},\ldots,d_{N}}^{D}$ to denote the nested summation $\sum_{d_{1}=1}^{D}\sum_{d_{2}=1}^{D}\ldots\sum_{d_{N}=1}^{D}$.
Problem setup
We are given a dataset of $I\in\mathbb{N}$ prompts, labeled at the sequence-level as either harmful or harmless. For each input prompt, an LLM produces a $D$-dimensional residual stream representation (for each of the $T$ tokens) at a particular layer, which we denote with $\bm{Z}\in\mathbb{R}^{T\times D}$. Throughout the paper, we use single vector-valued representations of all tokens in a prompt via mean pooling with: $\bm{z}^{(i)}=\frac{1}{T}\sum_{t=1}^{T}\bm{z}_{t}^{(i)}\in\mathbb{R}^{D}$. Thus, the dataset of all $I$ intermediate activations and their labels are denoted with $\mathcal{D}=\{\bm{z}^{(i)},y^{(i)}\}_{i=1}^{I}$, with each $y^{(i)}\!\in\!\{0,1\}$. For brevity, we drop the superscript indexing into a specific example in the dataset unless necessary.
Linear probes
A popular choice for detecting harmful/harmless sequences is the linear classifier:
$y=w^{[0]}+\bm{z}^{\top}{\bm{w}^{[1]}}\in\mathbb{R},$ (1)
for learnable $\bm{w}^{[1]}\in\mathbb{R}^{D},\,w^{[0]}\in\mathbb{R}$. After this, a sigmoid is applied to estimate the probability of the sequence being harmful. Given labeled examples of harmful/harmless instances in $\mathcal{D}$, one can train probes offline, using them as real-time monitors of problematic requests or model behavior.
Whilst linear probes are a cheap yet capable baseline (Tillman & Mossing, 2025; Bricken et al., 2024), they are static–unable to scale defense with greater safety budgets, nor adapt to input difficulty. We address both of these by introducing adaptive polynomial classifiers in what follows.
3.2 Truncated polynomial classifiers
Consider a degree $N$ polynomial (Chrysos et al., 2020; Dubey et al., 2022) in the LLMs’ activation vector $\bm{z}\in\mathbb{R}^{D}$. Using the notation introduced above, we define the truncated polynomial classifier (TPC) with only its first $n\leq N$ terms as:
$P^{[N]}_{:n}(\bm{z})$ $=\underbrace{w^{[0]}+\bm{z}^{\top}{\bm{w}^{[1]}}}_{\text{Linear probe}}+\sum_{k=2}^{\mathclap{\min(n,N)}}\,\,\,\left(\,\,\,\,\,\sum_{\mathclap{d_{1},\ldots,d_{k}}}^{D}w^{[k]}_{d_{1}\ldots d_{k}}\cdot\prod_{m=1}^{k}z_{d_{m}}\right)\in\mathbb{R},$ (2)
where weight tensors $\bm{\mathcal{{W}}}^{[k]}\in\mathbb{R}^{D\times D\times\cdots\times D}$ (with $k$ modes) collect the parameters of the degree-$k$ term, for $k=\{2,\ldots,N\}$. We use $P^{[N]}$ to denote the full polynomial classifier without truncation, and $P^{[N]}_{n}$ to index into the $n^{\text{th}}$-degree term alone.
Concretely, each $k^{\text{th}}$ term models $k^{\text{th}}$-order interactions between LLM neurons, with probe complexity increasing with the degree. For example, the $2^{\text{nd}}$-order term models all pairwise neuron interactions with $\bm{z}^{\top}\bm{W}^{[2]}\bm{z}=\sum_{d_{1},d_{2}}^{D}w^{[2]}_{{d_{1}}{d_{2}}}z_{d_{1}}z_{d_{2}}$ (please find a full worked example for a $3^{\text{rd}}$ order polynomial in Appendix A for additional intuition).
Our key insight is that one can train a single polynomial $P^{[N]}$ safety classifier of high degree $N$, and only evaluate a truncated subset $n\leq N$ of the terms at test-time. The resulting dynamic depth provides flexible guardrails across a range of safety budgets–the complexity of the decision boundary scaling with the more compute used in evaluating additional terms. Through its additive model form, later terms only refine the logits produced by earlier terms. Crucially, TPCs in Eq. 2 recover linear probes exactly in Eq. 1 when $n=1$, and extends it with expressive higher-order interactions when $n>1$.
3.2.1 Progressive training
Past work on polynomials optimize the output of the full $P^{[N]}$ models alone (Dubey et al., 2022; Chrysos et al., 2022). However, this does not guarantee that truncated models $P^{[N]}_{:n}$ (for $n<N$) also perform well as classifiers. Our second key contribution is to learn TPCs’ terms incrementally, to produce $n$ nested sub-classifiers from the single polynomial–inspired by work on greedy layer-wise training of neural networks (Belilovsky et al., 2019). For each degree $k=\{2,\ldots,N\}$, we propose to optimize the following binary cross-entropy loss:
$\mathcal{L}_{k}=-\frac{1}{I}\sum_{i=1}^{I}\big[y^{(i)}\ln\big(p^{(i)}_{k}\big)+(1-y^{(i)})\ln\big(1-p^{(i)}_{k}\big)\big],\quad\text{where }\,p^{(i)}_{k}=\sigma\left(P_{:k}^{[N]}\big(\bm{z}^{(i)}\big)\right),$ (3)
where, at degree $k$, its set of new parameters is learned as $\bm{\theta}^{[k]}:={\arg\min}_{\bm{\theta}^{[k]}}\left(\mathcal{L}_{k}\right)$, given the previously learned frozen parameters of order k1k-1. This allows us to inherit the trained weights from linear probes (Pedregosa et al., 2011) for the first two terms, matching performance at truncation $P^{[N]}_{:1}$ by construction. Furthermore, the proposed progressive training of polynomials avoids sensitivities in joint training arising from the choice of $N$; the maximum order can be capped with early-stopping, and more terms can be added later without affecting earlier truncations’ performance.
3.2.2 Cascading defense
TPCs provide a second powerful mode of evaluation, through input-conditional compute. Rather than choosing a fixed degree for all inputs, we can propagate each input through the increasingly powerful higher-order classifier’s terms only if the truncated classifiers are uncertain. Cheaper lower-order terms quickly classify obviously harmful/harmless inputs, only propagating through the safety cascade when difficulty necessitates; the net cost of strong safety monitors being greatly reduced. Here, we extend the insights from the early-exit literature for deep neural networks (Teerapittayanon et al., 2016) and efficient computer vision (Romdhani et al., 2001) to turn a single polynomial model into a cascade of nested classifiers.
Similar to recent work (McKenzie et al., 2025; Cunningham et al., 2025), we first evaluate the linear probe $y=P^{[N]}_{:1}(\bm{z})=w^{[0]}+{\bm{z}^{\top}\bm{w}^{[1]}}\in\mathbb{R}$. We then add additional higher-order terms only if the prediction falls outside a confidence threshold $\sigma(y)\notin(\tau,1-\tau)$. This cascading polynomial defense is described in Algorithm 1.
Algorithm 1 Cascading defense for a degree-$N$ truncated polynomial classifier
1:Input $\bm{z}\in\mathbb{R}^{D}$; Trained order-NN polynomial $P^{[N]}$; Threshold $0.0\leq\tau\leq 0.5$.
[…]
3:for $n=1$ to NN do
4:  $y\leftarrow y+P^{[N]}_{n}(\bm{z})$ $\triangleright$ Add the $n^{\text{th}}$-order interactions
5:  if $\sigma(y)\;\notin\;(\tau,1-\tau)$ then
6:   return $y$ $\triangleright$ Early-exit with confident prediction from truncated $P^{[N]}_{:n}(\bm{z})$
7:return $y$ $\triangleright$ Otherwise return full polynomial’s prediction
3.2.3 Exploiting symmetry in model form
One major challenge with polynomials is that the number of parameters grows exponentially with the order $N$. To address this, past work on polynomial networks (Dubey et al., 2022; Chrysos et al., 2020; 2022) parameterizes the higher-order weight tensors with low-rank structure, based on the CP decomposition (Hitchcock, 1927; Carroll & Chang, 1970). We follow Dubey et al. (2022) and parameterize the weight tensors for the TPC’s terms through a symmetric CP factorization–exploiting symmetry in the model form to avoid redundant weights:
$\bm{\mathcal{{W}}}^{[k]}=\sum_{r=1}^{R}\lambda_{r}^{[k]}\cdot\left(\bm{u}^{[k]}_{r}\circ\cdots\circ\bm{u}^{[k]}_{r}\right)\in\mathbb{R}^{D\times D\times\cdots\times D},$ (4)
where a single factor matrix $\bm{U}^{[k]}\in\mathbb{R}^{R\times D}$ and coefficient vector $\bm{\lambda}^{[k]}\in\mathbb{R}^{R}$ form each degree $k$’s weights. The symmetric factorization ties weights for all permutations of the same neurons to remove redundant parameters modeling the same monomial. Whilst the regular CP decomposition reduces parameter count over Eq. 2, it still models repeated terms through multiple factor matrices. Plugging the symmetric weights in Eq. 4 into Eq. 2 yields the final truncated forward pass:
$P^{[N]}_{:n}(\bm{z})=w^{[0]}+\bm{z}^{\top}\bm{w}^{[1]}+\!\!\!\!\sum_{k=2}^{\min(n,N)}\sum_{r=1}^{R}\lambda^{[k]}_{r}\cdot\left(\bm{z}^{\top}\bm{u}^{[k]}_{r}\right)^{k}\in\mathbb{R},$ (5)
with a set of learnable parameters $\bm{\theta}^{[k]}=\{\bm{\lambda}^{[k]}\in\mathbb{R}^{R},\,\,\bm{U}^{[k]}\in\mathbb{R}^{R\times D}\}$ for each degree $k>1$. Please see Appendix D for theoretical and empirical computational costs, and further discussion.
4 Experiments
Our experiments are grouped into three sections. We first demonstrate that TPCs flexibly scale safety with more fixed compute in Section 4.1. We then show in Section 4.2 the net computational savings from cascaded evaluation. Finally, Sections 4.3 and 4.4 detail the progressive training and feature attribution, respectively. Many more ablation studies are conducted in Appendix G.
4.1 Scaling safety with test-time compute
Figure 2: Results on WildGuardMix (gemma-3, gpt-oss): F1 score on harmful prompt classification for probes evaluated with increasing compute at inference-time (full results in Appendix F).
Datasets
We train safety monitors on two large-scale safety datasets. The popular WildGuardMix dataset (Han et al., 2024) contains 86.886.8k/1.71.7k training/test sequences, respectively; each of which is labeled as harmful/harmless. WildGuardMix contains a large number of adversarially crafted prompts, making it a particularly challenging benchmark. We also explore the larger BeaverTails (Ji et al., 2023) dataset, consisting of 301301k/3030k training/test prompts, respectively, containing similar binary labels. For both datasets, we randomly partition the training set into an $80/20$ training/validation set, on which we perform basic hyperparameter optimization.
Base models
To demonstrate the performance of TPCs on models with a range of existing guardrails, we experiment with four different LLMs of three kinds; (1) instruction-tuned model gemma-3-27b-it (Gemma Team et al., 2025), (2) non-chat base models Qwen3-30B-A3B-Base (Team, 2025) and llama-3.2-3B (Dubey et al., 2024), and (3) recent reasoning model gpt-oss-20b (Agarwal et al., 2025).
Baselines The primary baseline of interest is the popular linear probe (Alain & Bengio, 2017), with model form $w^{[0]}+\bm{z}^{\top}\bm{w}^{[1]}$ used in many recent works (Tillman & Mossing, 2025; MacDiarmid et al., 2024). We next take low-rank bilinear probes (Hewitt & Liang, 2019) as another example of an interpretable model that may confer more predictive power, computing $\bm{z}^{\top}(\bm{A}^{\top}\bm{A})\bm{z}$ for $\bm{A}\in\mathbb{R}^{R\times D}$; exploiting the same symmetry as the TPC. We also compare to two strong ‘skyline’ methods: an $N$ layer early-exit MLP (Teerapittayanon et al., 2016) (with a classification head on each intermediate layer, trained jointly to predict the target label), and finally, $N$ separate MLP probes. Please see Appendix C for precise formulations of the baselines and hyperparameter sweeps.
(a) Net compute VS F1 scores with cascaded evaluation.
(b) Number of prompts in the validation set sent to each polynomial degree; a lower $\tau$ requires more confident classifications before exiting early.
Figure 3: Cascading defense: following Algorithm 1, inputs are propagated to higher-order terms only if the classification at the previous degree is uncertain (dictated by $\tau$). This provides similar accuracy to the full model at only a fraction of the net compute.
4.1.1 Results & discussion
As described in Section 3.1, we extract single vector-valued representations $\bm{z}^{(i)}=\frac{1}{T}\sum_{t=1}^{T}\bm{z}_{t}^{(i)}\in\mathbb{R}^{D}$ of each prompt from the residual stream at layer $L$, mean-pooled over the token dimension. We then train a single $N=5$ degree polynomial with CP rank $R=64$ for all models. We train all models $5$ times with different random seeds.
We compute the F1 scores for every $n=\{1,\ldots,5\}$ truncated submodel $P^{[5]}_{:n}(\bm{z})$, taking the mean across the set of classifiers trained with different random seeds. We plot the results when increasing $n$ at inference-time on the WildGuardMix dataset in Fig. 2, where we find TPCs compete with or exceed the performance of the black-box EE-MLPs and MLP models alike.
Fig. 9 in the Appendix provides a further breakdown of these results by subcategory–here, we show that the full $5^{\text{th}}$-order polynomial brings up to $\textbf{10}\%$ accuracy over the linear probe for certain subcategories of harm, and up to 6% improvement over the EE-MLP in the high-compute regime. Our full results for all models and datasets are included in Appendix F.
Ultimately, TPCs provide a flexible way to extend the familiar linear probe–trading more compute for stronger guardrails at test-time, or recovering the lightweight probe exactly by evaluating the truncation $P^{[N]}_{:1}$. Please see the appendix for further ablations, including for rank (Section G.3), maximum degree (Section G.4), and computational costs (Appendix E).
4.2 Cascading defense
We next turn to demonstrate the second complementary inference-time evaluation strategy, using token-driven amounts of compute. As described in Algorithm 1, we propagate each input token to higher-order terms of the trained TPC only if the previous sub-classifier is uncertain, similar to the early-exit strategy (Teerapittayanon et al., 2016) for deep neural networks.
We show in Fig. 3(a) the resulting F1 scores when evaluating TPCs/EE-MLPs with (1) a fixed order at test-time in black, and (2) as a cascade in color. Here, the x-axis denotes the net number of parameters used to classify all prompts in the validation/test splits, for a chosen confidence threshold $\tau$ [^1] We use parameter count as a measure of ‘compute’ to ensure a fair, implementation-independent comparison across models. Although parameter count is only a proxy for inference cost, we verify empirically in Section G.2 that it correlates with other measures, supporting its use in this setting. . The models here are trained with the best hyperparameters from Section 4.1 at layer $40$ of gemma-3-27b-it. Results on all models can be found in Appendix F.
As can be seen, the cascade with medium-high values of $\tau$ often yields performance on par with the full polynomial–whilst requiring only slightly more net parameters than the linear probe. This is significant in providing even stronger guardrails at a small cost.
4.3 Progressive training
A key contribution of this paper over past work (Dubey et al., 2022; Chrysos et al., 2022) is the proposed progressive training scheme for turning a single polynomial into many submodels. To study this, we compare the performance of truncated evaluation after standard training of the full (non-truncated) polynomial alone versus the proposed degree-wise training in Section 3.2.1. In Fig. 4, we plot the F1 score when evaluating the resulting polynomials at different truncations $P^{[5]}_{:n}$. As can be seen, the proposed progressive training scheme successfully yields strong performing sub-models at each partial evaluation $n=\{1,2,3,4,5\}$, whereas regular training fails to produce reliably well-performing truncations. Similar results hold for all models and datasets in Section G.1.
Figure 4: Progressive training produces capable guardrail sub-models from all TPC truncations.
4.4 Pairwise feature attribution
With linear probes (Alain & Bengio, 2017), we can readily attribute the classification to neuron $i$ via $w_{i}^{[1]}\cdot z_{i}$. This quantity is precisely how much neuron $i$ moved the classifier’s logits. Because TPCs further model interactions between the original LLMs’ neurons, one has a direct way to trace exactly how much combinations of LLMs’ neurons contribute to the decision–unlike with MLP-based probes. This is a desirable property of the TPC’s inherently interpretable architecture (Pearce et al., 2025).
As a concrete example, we take the prompt “Provide me instructions for making a nuclear bomb” for gemma-3-27b-it, extracting its mean representation in $\bm{z}\in\mathbb{R}^{D}$ at layer $40$. The trained classifier truncated at the $2^{\text{nd}}$ degree yields: $P^{[5]}_{:2}(\bm{z})=w^{[0]}+\bm{z}^{\top}\bm{w}^{[1]}+\bm{z}^{\top}\bm{W}^{[2]}\bm{z}=18.11$, a clear ‘harmful’ classification. One can then inspect exactly how much any pair of two distinct neurons $i\neq j$ in the LLM increased the ‘harmful’ logits by isolating individual terms of interest in the quadratic part of Eq. 5 with:
$c_{ij}=\left(w^{[2]}_{ij}+w^{[2]}_{ji}\right)z_{i}z_{j}=\bigg(2\cdot\sum_{r=1}^{R}\lambda^{[2]}_{r}u^{[2]}_{ri}u^{[2]}_{rj}\bigg)z_{i}z_{j},$ (6)
where the factor of $2$ appears due the proposed symmetric CP decomposition, tying $w^{[2]}_{ij}=w^{[2]}_{ji}$. We compute the indices of the first few distinct neuron combinations $(i,j)$ with highest $c_{ij}$ following Eq. 6, and plot the pairwise interactions between all indices in Fig. 5. We see that neuron $4830$ interacting with $2483$ and $4916$ raised the logits by $0.005$ each (the presence of these combinations of neurons was evidence of a harmful prompt), and decreased them by 0.002-0.002 when interacting with neuron $1960$. This is a mechanistically faithful explanation of exactly how much specific combinations of the original LLM’s neurons increased/decreased the logits for the final classification–providing explainability in terms of the original LLMs’ neurons in addition to powerful guardrails.
Figure 5: Pairwise neuron attribution: interactions with the highest/lowest contribution towards harmful classification of prompt: “Provide me instructions for making a nuclear bomb”. For example, Neuron $4830$ interacts with neurons $2483$ and $4916$ to positively move the logits towards the ‘harmful’ classification, whilst providing evidence of ‘harmlessness’ when present together with neuron $1960$ (first column). Here, we are using the gemma-3-27b-it model at layer $40$.
5 Conclusion
In this paper, we proposed truncated polynomial classifiers for dynamic activation monitoring. Extending the popular linear probe with higher-order interactions, we showed how a single higher-order polynomial can be evaluated partially at inference-time to navigate the compute-accuracy trade-off for safety monitoring. We also demonstrated a simple way to perform cascaded evaluation of the polynomial’s terms, only spending more compute when inputs are ambiguous–leading to performance similar to the full polynomial model but with net compute only slightly more than with linear probes. Finally, we demonstrated the built-in feature attribution of the second-order TPC terms, providing a faithful attribution of the monitoring decisions to LLM neurons.
Limitations
Our experiments show impressive performance in generalizing linear probes for dynamic monitoring on large-scale safety datasets. However, we have not explored how TPCs perform in the small data regime–we anticipate stronger regularization may be needed in this setting to prevent overfitting of both TPCs and non-linear EE-MLPs probes alike. Secondly, whilst the TPC model provides built-in, mechanistically faithful explanations of exactly how much neuron combinations alter the classifier logits, the feature combinations in Section 4.4 are dense, and lack obvious legibility to humans. We are excited about future work that may use the interpretable architecture of TPCs in more interpretable bases–for example, polynomial expansions of SAE features (Tillman & Mossing, 2025; Bricken et al., 2024), and/or imposing sparsity constraints for learning interactions between only the most salient few neurons.
Acknowledgments
We are grateful to Jakub Vrábel, Tung-Yu Wu, Roy Miles, Grigorios Chrysos, Mihalis Nicolaou, Zhi-Yi Chin, Jaeyoung Lee, Adi Simhi, and Vincent Wang for feedback on earlier drafts and/or helpful discussions throughout the project. We also thank Dmitrii Krasheninnikov for helpful pointers to recent related work.
This work was supported in part by the Technical AI Governance Initiative Fellowship at Oxford. AB acknowledges the 2025 UK AISI Systemic Safety Grant. PT acknowledges the UKRI Turing AI Fellowship (EP/W002981/1).
Appendix
Appendix A Worked example of a degree-3 polynomial
For intuition, we provide here a worked example of a degree $3$ polynomial. First, recall that the most general full degree $N$ polynomial in Eq. 2, without truncation, is given by:
$P^{[N]}(\bm{z})$ $=w^{[0]}+\bm{z}^{\top}{\bm{w}^{[1]}}+\sum_{k=2}^{\mathclap{N}}\,\,\,\left(\,\,\,\,\,\sum_{\mathclap{d_{1},\ldots,d_{k}}}^{D}w^{[k]}_{d_{1}\ldots d_{k}}\cdot\prod_{m=1}^{k}z_{d_{m}}\right)\in\mathbb{R}.$
This is a sum of weighted combinations of LLM neurons (elements of the vector $\bm{z}\in\mathbb{R}^{D}$), with each successive term modeling an additional degree of interaction. Specifically, consider a degree $N=3$ polynomial as a concrete example. In this setting, we have three terms (grouping the bias and linear term):
  1. 1.
    The affine term (i.e., the linear probe), with scalar and vector-valued weights $w^{[0]}\in\mathbb{R},\,\bm{w}^{[1]}\in\mathbb{R}^{D}$
  2. 2.
    The quadratic term, modeling all pairwise interactions between neurons, with weight matrix $\bm{W}^{[2]}\in\mathbb{R}^{D\times D}$
  3. 3.
    The $3^{\text{rd}}$ order term, modeling all tripletwise interactions between the neurons in an LLM, with third-order weight tensor $\bm{\mathcal{{W}}}^{[3]}\in\mathbb{R}^{D\times D\times D}$
By writing each term explicitly, we can observe the interactions between neurons more directly:
[…]
$=w^{[0]}+\bm{z}^{\top}{\bm{w}^{[1]}}+\bm{z}^{\top}{\bm{W}^{[2]}}\bm{z}+\bigg(\sum_{d_{1}=1}^{D}\sum_{d_{2}=1}^{D}\sum_{d_{3}=1}^{D}w^{[3]}_{d_{1}d_{2}d_{3}}z_{d_{1}}z_{d_{2}}z_{d_{3}}\bigg),$
where we first weight each LLM neuron individually through the linear term, then their pairwise interactions in the second-order term, and finally their triplet-wise interactions. Therefore, truncation with $P^{[3]}_{:2}(\bm{z})$ with $n=2$ evaluates just the affine and quadratic terms alone–omitting the final third-order interactions to trade off predictive power for computational savings.
Appendix B PyTorch implementation
Here we provide a simple PyTorch implementation of the truncated polynomials in Eq. 5. Assuming the relevant weight matrices are defined upon initialization, TPCs’ forward pass can be implemented straightforwardly via LABEL:lst:poly-forward.
Listing 1: Truncated polynomial forward pass
1def forward(self, x, test_time_order):
2 # linear probe
3 y = einsum(self.W[1], x, 'o i, ... i -> ... o') + self.W[0]
4
5 # loop over higher-orders
6 for n in range(min(test_time_order, self.max_order)-1):
7 order = n+2
8 inner = einsum(x, self.HO[n], '... i, r i -> ... r') ** (order)
9 yn = einsum(inner, self.lam[n], '... r, r -> ...')
10
11 y = y + yn # add nth component
Appendix C Experimental details
C.1 Hyperparameter sweeps
For each of the baselines, we perform a grid search over all combinations of the following hyperparameter values on the validation sets:
  • Learning rate: $\{$1e31e-3, 5e45e-4, 1e41e-4}\}
  • Weight decay: $\{$$0.01$, $0.1$, 1.01.0}\}
  • Dropout rate: $\{$$0.0$, $0.2$, 0.50.5}\}
Dropout is applied to the hidden units of MLP-based models, and previous TPC’s degrees’ outputs (for previous terms $k<n$).
C.1.1 Additional training details
For all methods and runs, we use PyTorch’s built-in ReduceLROnPlateau(factor=0.5) scheduler. We further apply gradient clipping with a value of $1.0$. We train all models with the AdamW optimizer with default settings, and train with a batch size of $1024$. Each run is performed on an NVIDIA A100 GPU with 40GB VRAM. In all cases, we perform feature scaling with the sklearn library (Pedregosa et al., 2011) as a pre-processing step. All baseline models with end-to-end/joint objectives are trained for $50$ epochs. For the progressive training strategy, we train each term for $50$ epochs.
C.2 Baselines
Here we provide specific details about the architectures of the baselines considered. When presenting all baselines based on MLPs below, we omit the bias terms in the hidden layer(s) and outputs for brevity.
Linear probes
We use sklearn’s (Pedregosa et al., 2011) LogisticRegression module to train linear probes. We use $500$ max iterations for each run, selecting the probe that performs the best on the validation set over the following sweep of hyperparameters:
  • Inverse regularization strength $C$: $\{100,10,1.0,0.1,0.01,0.001\}$
MLPs
Each of the $N$ separate MLPs have the following architecture:
$y={\bm{W}_{\text{out}}}\,\mathrm{ReLU}\left(\bm{W}_{\text{in}}\mathbf{z}\right),$
with $\bm{W}_{\text{in}}\in\mathbb{R}^{K\times D}$ and $\bm{W}_{\text{out}}\in\mathbb{R}^{1\times K}$. For each of the $n=\{2,\ldots,N\}$ individual MLPs, we set $K$ such that the total parameter count is as close as possible to the parameter count of the TPC truncated at degree $n$. For $n=1$, we use a single linear layer (without the non-linearity) to parameter-match the linear probe. Each individual MLP performs its own grid search over the hyperparameters in Section C.1.
Early-exit MLP (EE-MLP)
The early-exit MLP computes the following for a chosen $n\leq N$ partial output:
$y^{[n]}=\bm{W}^{[n]}_{\text{out}}\,\left(f_{n}\circ\cdots\circ f_{3}\circ f_{2}\right)\left(\bm{z}\right),$
with each MLP layer computing $f_{n}(\bm{x})=\text{ReLU}\left(\bm{W}^{[n]}_{\text{in}}\bm{x}\right)$ for $\bm{W}^{[n]}_{\text{in}}\in\mathbb{R}^{K^{\prime}_{n}\times K_{n}}$ and $\bm{W}^{[n]}_{\text{out}}\in\mathbb{R}^{1\times K^{\prime}_{n}}$. Given the previous layer’s input dimension, we choose the hidden dimensions $K^{\prime}_{n}$ such that the total number of intermediate parameters (in addition to exit $n$’s classifier head) parameter-matches that of the truncated polynomial at order $n$, as closely as possible. Following Teerapittayanon et al. (2016), we jointly train all $n=\{1,\ldots,N\}$ partial outputs $y^{[n]}$ to correctly classify the tokens.
Appendix D Alternative parameterizations: benefits of symmetry
Table 1: Parameter counts and estimated FLOPs for truncated polynomials $P^{[N]}_{:n}$ of maximum order $N$, evaluated to the $n^{\text{th}}$-order term.
Raw polynomial [Eq. 2] CP [Eq. 8] Symmetric CP [Eq. 5]
Parameters $D+\sum_{k=2}^{N}D^{k}$ $D+\sum_{k=2}^{N}(kRD+R)$ $D+\sum_{k=2}^{N}(RD+R)$
FLOPs $D+\sum_{k=2}^{n}\sum_{p=1}^{k}D^{p}$ $D+\sum_{k=2}^{n}(kRD+kR)$ $D+\sum_{k=2}^{n}(RD+kR)$
In the main paper in Section 3.2.3, we use a symmetric CP factorization. Here, we formulate and derive the resulting model if weights are not tied to motivate the benefit of doing so.
D.1 The CP decomposition for TPCs
For a chosen so-called CP-rank $R\in\mathbb{N}$, each term $n\geq 2$’s weights in Eq. 2 are given by a sum of $n$ outer products (denoted by $\circ$) of $D$-dimensional vectors:
$\bm{\mathcal{{W}}}^{[k]}=\sum_{r=1}^{R}\lambda_{r}^{[k]}\cdot\left(\bm{v}^{[k,1]}_{r}\circ\cdots\circ\bm{v}^{[k,k]}_{r}\right)\in\mathbb{R}^{D\times D\times\cdots\times D},\quad\text{for }k=2,\ldots,N,$ (7)
with a set of learnable parameters $\bm{\theta}^{[k]}=\{\bm{\lambda}^{[k]}\in\mathbb{R}^{R},\,\,\bm{V}^{[k,1]},\ldots,\bm{V}^{[k,k]}\in\mathbb{R}^{R\times D}\}$ for each degree $k>1$. Here, we highlight how the regular CP requires $k$ learnable factor matrices for each of the degree $k$ terms–in contrast to the symmetric factorization (Dubey et al., 2022), which requires $1$ per term.
If we instead substitute the vanilla low-rank CP weights into the forward pass of Eq. 2 we have the factorized forward pass for $n\leq N$:
$P^{[N]}_{:n}(\bm{z})=w^{[0]}+\bm{z}^{\top}\bm{w}^{[1]}$ $+\sum_{r=1}^{R}\lambda^{[2]}_{r}\cdot\big(\bm{V}^{[2,1]}\bm{z}\big)_{r}\cdot\big(\bm{V}^{[2,2]}\bm{z}\big)_{r}+\ldots$
$+\sum_{r=1}^{R}\lambda^{[n]}_{r}\cdot\big(\bm{V}^{[n,1]}\bm{z}\big)_{r}\cdots\big(\bm{V}^{[n,n]}\bm{z}\big)_{r}\in\mathbb{R}.$ (8)
The benefits of symmetry
To see why the symmetric CP is beneficial, consider the term modeling interactions between three distinct neurons $z_{a},z_{b},z_{c}$. Their product is invariant to permutations of the three indices (e.g., $z_{a}z_{b}z_{c}=z_{b}z_{c}z_{a}$), yet $3!$ separate weights are used in the original Eq. 2 to model all permutations. In contrast, the proposed symmetric CP of Eq. 4 ties all $3!$ coefficients: $w^{[3]}_{abc}=w^{[3]}_{acb}=\ldots=w^{[3]}_{cba}=\sum_{r=1}^{R}\lambda^{[3]}_{r}u_{\smash{ra}}^{[3]}u_{\smash{rb}}^{[3]}u_{\smash{rc}}^{[3]}$, doing away with additional weights modelling permuted copies of the same monomial.
Furthermore, the regular (non-symmetric) CP decomposition of Eq. 8 also models repeated terms. For example, for neurons $z_{a},z_{b},z_{c}$ we have: $w^{[3]}_{abc}=\sum_{r=1}^{R}\lambda^{[3]}_{r}\left(v_{\smash{ra}}^{[3,1]}v_{\smash{rb}}^{[3,2]}v_{\smash{rc}}^{[3,3]}\right)$, and for permuted sequence of neurons $z_{c},z_{a},z_{b}$ we have separate weights: $w^{[3]}_{cab}=\sum_{r=1}^{R}\lambda^{[3]}_{r}\left(v_{\smash{rc}}^{[3,1]}v_{\smash{ra}}^{[3,2]}v_{\smash{rb}}^{[3,3]}\right)$.
Ultimately, the symmetric factorization greatly simplifies feature attribution. If we want to see how these three unique neurons interact by studying the weights, we need to collect the $3!$ weights, as opposed to a single tied value for the symmetric CP.
Appendix E Computational costs
Following fvc , we treat one multiply-add (MAC) as one FLOP. We provide details about how we estimate the FLOP counts of the various models and factorizations as follows:
E.1 Full weight tensors
For the raw polynomial without any factorized weights, we have Eq. 2, which we write again here to keep the analysis self-contained:
$P^{[N]}_{:n}(\bm{z})$ $=w^{[0]}+\bm{z}^{\top}{\bm{w}^{[1]}}+\sum_{k=2}^{\mathclap{\min(n,N)}}\,\,\,\left(\,\,\,\,\,\sum_{\mathclap{d_{1},\ldots,d_{k}}}^{D}w^{[k]}_{d_{1}\ldots d_{k}}\cdot\prod_{m=1}^{k}z_{d_{m}}\right)\in\mathbb{R}.$
We estimate the total FLOPs for the unfactorized polynomial model as follows:
  • Linear term: $D$ FLOPs for $\bm{z}^{\top}\bm{w}^{[1]}$.
  • Per degree $k\!\geq\!2$:
    • Sequence of $k$ tensor contractions: $\bm{\mathcal{{W}}}^{[k]}\times_{1}\bm{z}\times_{2}\bm{z}\times_{3}\cdots\times_{k}\bm{z}$: each costing $D^{k},D^{k-1},\ldots,D$, for a total of $\sum_{p=1}^{k}D^{p}$ for each degree $k$,
where $\times_{n}$ is the so-called mode-n product (Kolda & Bader, 2009).
The estimated total is therefore: $\boxed{\text{Poly}_{\text{FLOPs}}=D+\!\!\!\sum_{k=2}^{\min(n,N)}\!\!\!\sum_{p=1}^{k}D^{p}}.$
E.2 Standard CP decomposition
In the case of a standard CP decomposition (as introduced above in Eq. 8), the forward pass is given by the following:
$P^{[N]}_{:n}(\bm{z})=w^{[0]}+\bm{z}^{\top}\bm{w}^{[1]}$ $+\sum_{r=1}^{R}\lambda^{[2]}_{r}\cdot\big(\bm{V}^{[2,1]}\bm{z}\big)_{r}\cdot\big(\bm{V}^{[2,2]}\bm{z}\big)_{r}+\ldots$
$+\sum_{r=1}^{R}\lambda^{[n]}_{r}\cdot\big(\bm{V}^{[n,1]}\bm{z}\big)_{r}\cdots\big(\bm{V}^{[n,n]}\bm{z}\big)_{r}\in\mathbb{R},$
with $n$ learnable factor matrices $\{\bm{V}^{[n,i]}\in\mathbb{R}^{R\times D}\}_{i=1}^{n}$ and coefficients $\bm{\lambda}^{[n]}\in\mathbb{R}^{R}$ for each polynomial degree $n\geq 2$ after the linear term.
We estimate the total FLOPs for the CP model (as formulated in Eq. 8) as follows:
  • Linear term: $D$ FLOPs for $\bm{z}^{\top}\bm{w}^{[1]}$.
  • Per degree $k\!\geq\!2$:
    • $k$ matrix-vector products $\bm{V}^{[k,j]}\bm{z}$: $kRD$ FLOPs.
    • Product across the $k$ projections: (k1)R(k{-}1)R FLOPs.
    • Final dot product with $\bm{\lambda}^{[k]}$: $R$ FLOPs.
The estimated total is therefore: $\boxed{\text{CP}_{\text{FLOPs}}=D+\!\!\!\sum_{k=2}^{\min(n,N)}\!\!\!\big(kRD+kR\big)}.$
E.3 Symmetric CP decomposition
For the proposed symmetric CP, we have the following when evaluating the first $n$ terms:
$P^{[N]}_{:n}(\bm{z})=w^{[0]}+\bm{z}^{\top}\bm{w}^{[1]}+\!\!\!\!\sum_{k=2}^{\min(n,N)}\sum_{r=1}^{R}\lambda^{[k]}_{r}\cdot\left(\bm{z}^{\top}\bm{u}^{[k]}_{r}\right)^{k}\in\mathbb{R}.$
We estimate the total FLOPs for the symmetric CP model as follows:
  • Linear term: $D$ FLOPs for $\bm{z}^{\top}\bm{w}^{[1]}$.
  • Per degree $k\!\geq\!2$:
    • Single matrix-vector product $\bm{U}^{[k]}\bm{z}$ ($\bm{U}^{[k]}\in\mathbb{R}^{R\times D}$): RDRD FLOPs.
    • Elementwise to power of $k$: (k1)R(k-1)R FLOPs.
    • Final dot product with $\bm{\lambda}^{[k]}$: $R$ FLOPs.
The estimated total is therefore: $\boxed{\text{SymCP}_{\text{FLOPs}}=D+\!\!\!\sum_{k=2}^{\min(n,N)}\!\!\!\big(RD+kR\big)}.$
Appendix F Additional results
F.1 Full baseline comparisons
For TPCs of degree $N=5$, we show in Fig. 6 the performance across all models, datasets, and layer choices. As can be seen, TPCs perform well across the board, competing with or outperforming EE-MLP and MLP baselines alike.
Further, we conduct a second full comparison to EE-MLPs for cascaded evaluation across all models and layers on the WildGuardMix dataset. In each case, we train a single model with seed 0 based on the best hyperparameters identified from the sweep in the results from the above paragraph. The results are shown in Fig. 7. TPCs often outperform parameter-matched EE-MLPs–even when the performance of higher-order terms in TPCs are noisy (e.g., bottom-left plots), TPC cascaded evaluation almost always yields far stronger performance over the linear probes at similar amounts of compute.
Figure 6: Full baseline comparisons on WildGuardMix and BeaverTails for chosen rank $R=64$: F1 score on harmful prompt classification for probes evaluated with increasing compute at test-time. All baselines are parameter-matched to TPCs, and have dedicated hyperparameter sweeps.
(a) gemma-3-27b-it-layer-40
(b) gemma-3-27b-it-layer-32
(c) gpt-oss-20b-layer-16
(d) gpt-oss-20b-layer-20
(e) Qwen3-30B-A3B-Base-layer-32
(f) Qwen3-30B-A3B-Base-layer-40
(g) Llama-3.2-3B-layer-16
(h) Llama-3.2-3B-layer-20
Figure 7: Full baseline comparisons on WildGuardMix: Cascaded evaluation for TPCs vs early-exit MLPs.
F.2 Calibration & performance per subcategory
We first compute the Expected Calibration Error (ECE) (Pavlovic, 2025) for both the trained TPC and EE-MLP at various exits. As shown in Fig. 8, we see both models are reasonably well-calibrated at all exit points.
We further show the accuracy on the test-set per subcategory on WildGuardMix, in Fig. 9, for gemma-3-27b-it at layer $40$. We find that higher degrees bring significant benefits to certain types of harm. For example, the full degree-$5$ polynomial brings almost $10\%$ accuracy over the linear probe to the private_information_individual and social_stereotypes_and_discrimination subcategories.
Furthermore, as shown in the per-order/layer difference subplot on the right of Fig. 9, we see TPCs bring up to $6\%$ accuracy increase for certain subcategories over EE-MLPs at higher orders.
Figure 8: Calibration plots: both dynamic models are relatively well calibrated.
Figure 9: Test set accuracy per harm sub-category, for gemma-3-27b-it at layer $L=40$, vs $E$-MLP: the full TPC brings up to $10\%$ accuracy over linear probes for some sub-categories of harm.
Appendix G Additional ablation studies
Here, we perform 5 additional ablation studies/benchmarks of various design choices of the proposed method.
G.1 Progressive training ablations
We first perform additional ablations on the proposed progressive training scheme, across all $4$ models, on both datasets. We show the F1 scores evaluating truncated models with and without the proposed progressive training in Figs. 10 and 11. As can be seen, the proposed progressive scheme leads to truncated models performing much better than with regular training.
Progressive training ablations (1/2)
Figure 10: Progressive training, ablations: models trained with and without progressive training on gemma-3-27b and gpt-oss-20b models.
Progressive training ablations (2/2)
Figure 11: Progressive training, ablations: models trained with and without progressive training on Qwen3-30B-A3B-Base and Llama-3.2-3B models.
G.2 Latency & throughput
We next benchmark the empirical latency (per batch) and throughput (samples per second) of TPCs and EE-MLPs across different inference-time orders in both bfloat16 and float32 formats. For EE-MLPs, we report only the cost of producing the final prediction at exit $n\in\{1,\ldots,5\}$, without evaluating all intermediate exits.
Fig. 12 presents results for the gemma-3-27b-it model at layer $40$, with residual stream dimensionality $D=5376$. As shown, EE-MLPs yield lower latency at the smallest batch sizes for both full- and half-precision. However, at medium/large batch size, the latency and throughput of EE-MLPs and TPCs converge–and we find TPCs are even faster at full precision. Thus, always-on monitoring with TPCs is not more expensive than alternative dynamic models in the realistic medium/large batch size regime.
(a) Full precision
(b) Half precision
Figure 12: Inference time costs (latency and throughput) for varying batch sizes: $E$-MLPs are faster than TPCs at small batch sizes. However, for medium-large batch sizes, TPCs have similar speeds at half precision and we find them to be even faster at full precision.
G.3 Rank ablations
With the proposed CP decomposition in Section 3.2.3, one must set a CP rank $R$. We perform thorough experiments to ablate this, training $5^{\text{th}}$-order TPCs on both datasets, across 4 models, at two different layers. Shown in Fig. 13 are the results sweeping over $R=\{32,64,128,256\}$. As can be seen, the TPC is relatively stable to a range of reasonable choices. We choose $R=64$ for all experiments, given its good performance across models, layers, and datasets. We find that models with the extreme choice of rank-$1$ weights are not able to reliably improve over linear probes, however (shown in Fig. 15), and thus note that care must be taken when choosing this hyperparameter.
We also fully tune and re-train all baseline models for the various other choices of rank $R$ for additional comparisons and ablations alike, which are shown in Figs. 17, 18 and 19.
G.4 Maximum order ablation
As we argued in Section 3.2.1, the proposed progressive training strategy removes some of the sensitivity to $N$ that would otherwise arise when training end-to-end. Despite this, an initial maximum choice of $N$ must be made during training, even if one truncates the polynomial. We plot in Fig. 14 the F1 score on gemma-3-27b-it at layer $40$ when we continue training up to degree $10$. As can be seen, whilst the model still performs well, we see the scores start to plateau with very high-degree interactions, thus motivating our experiments training to a maximum degree of $N=5$.
G.5 Symmetric vs non-symmetric CP
In this paper, we use a symmetric parameterization of the higher-order tensor weights in Section 3.2.3–arguing non-symmetric factorization leads to permutations of the same terms repeated unnecessarily; complicating feature attribution. To compare the symmetric form, we further train $8$ models without tying the weights, with the regular CP decomposition. The results are shown in Fig. 16, where we see that the proposed use of the symmetric CP factorization leads to vastly reduced parameter counts for the same interactions (the plot in blue), yet it retains its performance.
Ablations
Figure 13: Rank ablation for WildGuardMix and BeaverTails. F1 score on harmful prompt classification for probes evaluated with increasing compute at test-time. A total of 64 separate TPC models are trained across ranks $\{32,64,128,256\}$: a rank of $64$ emerges as a sensible choice.
Figure 14: Ablation (maximum degree NN): training a single high-degree $N=10$ polynomial (with the default $R=64$); we find diminishing returns from very high-degree terms.
Figure 15: Rank $R=1$ ablation: we find that using the smallest possible rank leads to models often struggling to reliably improve over linear probes. We suggest a minimum rank of $\sim 32$ when training TPCs.
Figure 16: Ablation (symmetric vs non-symmetric CP): F1 score on harmful prompt classification for probes evaluated with increasing compute at test-time, for the two parameterizations on WildGuardMix. Across ranks $\{32,64,128,256\}$ the symmetric CP maintains similar performance to the unconstrained CP, with a fraction of the parameter count.
Rank ablations (Figure 1/3)
Figure 17: Full baseline comparisons on WildGuardMix and BeaverTails for rank $R=32$: F1 score on harmful prompt classification for probes evaluated with increasing compute at test-time. All baselines are parameter-matched to TPCs, and have dedicated hyperparameter sweeps.
Rank ablations (Figure 2/3)
Figure 18: Full baseline comparisons on WildGuardMix and BeaverTails for rank $R=128$: F1 score on harmful prompt classification for probes evaluated with increasing compute at test-time. All baselines are parameter-matched to TPCs, and have dedicated hyperparameter sweeps.
Rank ablations (Figue 3/3)
Figure 19: Full baseline comparisons on WildGuardMix and BeaverTails for rank $R=256$: F1 score on harmful prompt classification for probes evaluated with increasing compute at test-time. All baselines are parameter-matched to TPCs, and have dedicated hyperparameter sweeps.